The speed and wavelength of coastal water waves are 4.0 m/s and 6.0 m respectively. When the waves enter a shallower region, the speed decreases to 1.7 m/s. What is the wavelength in the shallower region?


1
Expert's answer
2021-07-23T12:51:13-0400

The frequency of the waves is given as follows:


"f = \dfrac{v_1}{\lambda_1}"

where "v_1 = 4m/s, \lambda_1 = 6m". Assuming the frequency did not change, the wavelength in the shallower region is given as follows:


"\lambda_2 = \dfrac{v_2}{f} = \dfrac{v_2\lambda_1}{v_1}"

where "v_2 = 1.7m\/s". Thus, obtain:


"\lambda_2 = \dfrac{1.7\cdot 6}{4} \approx 2.6m"

Answer. 2.6 m.